OpenStudy (anonymous):

find the derivative of sqrt(3 x^2-2)

OpenStudy (anonymous):

\[\sqrt{3x^2 - 2}\] \[(3x^2 - 2)^{1/2}\] Chain rule. \[6x * 1/2 (3x^2 - 2) ^{-1/2}\] \[(3x)/(\sqrt{3x^2 - 2})\]

OpenStudy (anonymous):

jmtrapasso is incorrect.

OpenStudy (anonymous):

yes i see if im looking for an equation of the tangent line at 3 5 do i just plug in 5 to find f of x

OpenStudy (anonymous):

If you are looking for the equation of a tangent line at a point, plug in the x value of the point in f'(x) to get the slope at that point. Then, use point-slope formula. \[y-y _{1} = m(x-x _{1})\]

OpenStudy (anonymous):

i got it thanks